OpenStudy (anonymous):

22. Solve |x-1/5|=3. (1 point) x = –14/5 and x =16/5 x = 14/5 and x = –16/5 x = –14/5 and x = –16/5 x = 14/5 and x = 16/5

OpenStudy (anonymous):

x- 1/5 = 3 or x-1/5 = -3 x= 16/5 or x=-14/5